OpenStudy (anonymous):

solve for x. use substitution principle to check you answers. 1. |x-17|=25 2. |2x+3|=5 3. |7x-2|=12 4. |12+4x|=16 5. |3x-10|=17 6. |65+5x|=10 7. |17x-1|=0 8. |2/3x+4|=18

OpenStudy (anonymous):

@amonoconnor

OpenStudy (anonymous):

1.42 2. x=1 3.x=2

OpenStudy (stormoshawty):

1. x=42 2. x=1 3. x=2 4. x=1 5. x=9

OpenStudy (anonymous):

ok thank u , how bout 7,8, nd 9 ?

OpenStudy (stormoshawty):

I'm working on six but it's confusing. Give me a minute.

OpenStudy (stormoshawty):

I believe that 6 is x=-11

OpenStudy (anonymous):

ok

OpenStudy (ranga):

Whenever you have the absolute operator in an equation or inequality split the problem into two parts while getting rid of the absolute operator: |x-17|=25 means +(x - 17) = 25 or -(x - 17) = 25 x - 17 = 25. x = 25 + 17 = 42 -(x - 17) = 25 -x + 17 = 25. 17-25 = x or x = -8 Answer x = -8 or 42

OpenStudy (anonymous):

ive never takn algebra two so i know nothing

OpenStudy (ranga):

Each of the equations has two solutions. I just did the first one. You can put x = -8 or 42 in the first equation and you will see the equation is satisfied for both x values. You do similar thing for the remaining questions.

OpenStudy (ranga):

I will do the last one: |2/3x+4|=18 means +(2/3x + 4) = 18 or -(2/3x + 4) = 18 +(2/3x + 4) = 18 Add -4 to both sides 2/3x = 14 multiply both sides by 3 2x = 42 divide both sides by 2 x = 21 -(2/3x + 4) = 18 multiply both sides by -1 2/3x + 4 = -18 add -4 to both sides 2/3x = -22 multiply both sides by 3 2x = -66 divide both sides by 2 x = -33 Solution: x = 21 or -33

OpenStudy (anonymous):

would one the answers for number 2 , x=-4?

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

4. x=1 , x=-2 ?

OpenStudy (ranga):

x = -2 won't work for 4.

OpenStudy (anonymous):

oh , darnittt

OpenStudy (ranga):

|12+4x|=16 12 + 4x = 16 or -12 - 4x = 16 4x = 4 x = 1 -12 - 4x = 16 -12 - 16 = 4x -28 = 4x x = -7 x = -7 or 1

OpenStudy (anonymous):

i cant figure out 6 either

OpenStudy (ranga):

It is the same procedure. |65+5x|=10 means: +(65 + 5x) = 10 or -(65 + 5x) = 10 +(65 + 5x) = 10 add -65 to both sides 5x = 10 - 65 = -55 divide both sides by 5 x = -11 -(65 + 5x) = 10 -65 - 5x = 10 add 65 to both sides -5x = 10 + 65 = 75 divide both sides by -5 x = -15 x = -11 or -15

OpenStudy (anonymous):

im just havin problems finding the second number on 3 and 5 on 3 i got x=2 and on 5 i got x=9 just cant find the other ones . and i cant find either on 7

OpenStudy (ranga):

If you learn to do it properly once you can do all the others. The second one on 3 will be: -(7x-2) =12 I will make it easier. Multiply booth sides by -1 and tell me what you get.

OpenStudy (anonymous):

-7x-12=2?

OpenStudy (anonymous):

-7x=14?

OpenStudy (anonymous):

x=21

OpenStudy (ranga):

you have to be careful with the signs. I have changed the method slightly to hopefully make it easier. -(7x-2) =12 multiply both sides by -1 7x - 2 = -12 add 2 to both sides 7x = -12 + 2 = -10 divide both sides by 7 x = -10/7

OpenStudy (ranga):

Q5 is |3x-10|=17 you are having trouble with the second part or the negative part. -(3x - 10) = 17 multiply both sides by -1 and tell me what you get.

OpenStudy (anonymous):

-1.4?

OpenStudy (anonymous):

for number3

OpenStudy (ranga):

you can leave it as a fraction unless the answers are in decimal or they specifically tell you to answer in decimals.

OpenStudy (anonymous):

my answer was in decimals

OpenStudy (ranga):

then -1.4

OpenStudy (anonymous):

-|3x-10|=17 3x-10=17 +10 +10 3x=27 x=9

OpenStudy (ranga):

ok, you got the positive part. try the negative part next.

OpenStudy (anonymous):

thats what i just did ?

OpenStudy (ranga):

you had the negative sign in the first line but it disappeared from line 2 onwards.

OpenStudy (anonymous):

3x-10=-17 3x=-7 x=-2.3

OpenStudy (ranga):

Yes!

OpenStudy (anonymous):

so for number 7: positive: |17x-1|=0 17x-1=0 17x=1 x=17?

OpenStudy (ranga):

17x = 1 divide both sides by 17

OpenStudy (anonymous):

1x=0.06

OpenStudy (ranga):

yeah, x = 1/17 = 0.06 do the negative part next.

OpenStudy (ranga):

because the right side is zero, you will get the same answer for negative as you got for positive.

OpenStudy (anonymous):

-|17x-1|=0 17x-1=0 17x=-1?

OpenStudy (ranga):

By the way, when you put the negative sign the absolute bar should not be there on line 1. It should be parenthesis. We split the equation into two parts to get rid of the absolute bars and so no need to use them anymore. -(17x - 1) = 0 you got to watch the sign -17x + 1 = 0 add 17x to both sides 1 = 17x divide by 17 1/17 = x same as for positive.

OpenStudy (anonymous):

so both 0.06 ?

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

thank you so much (:

OpenStudy (ranga):

So for the last one you will say x = 0.06 just once. All the others had two solutions but this one just had one. You are welcome.